# SAFE: Multitask Failure Detection for Vision-Language-Action Models

### v1

@@ 4.3 Threshold Selection by Conformal Prediction
When the predicted failure score $s_{t}$ exceeds the time-varying threshold $\delta_{t}$, we raise a failure flag. To determine $\delta_{t}$ in a principled way, we adopt the functional conformal prediction (CP) framework [22]. Functional CP constructs a time-varying prediction band by leveraging the distribution of $s_{t}$ observed in successful rollouts within a calibration set. Under the exchangeability assumption [58] and given a user-specified significance level $\alpha$, the CP band guarantees that, for a new successful rollout, its $s_{t}$ will lie within this band at all times $t$ with probability $1-\alpha$. Conversely, if the score of a test rollout exits the band at time $t$, we can declare a failure with nominal confidence $1-\alpha$.
Formally, given a time series of any scalar score $s_{t}$ and a user-specified significance level $\alpha\in(0,1)$, functional CP gives a distribution-free prediction band $C_{\alpha}$. Following [8], we adopt the one-sided time-varying CP band formulation, where $C_{\alpha}$ is a set of intervals $\{[\text{lower}_{t},\text{upper}_{t}]:t=1,\dots,T\}$, where $\text{lower}_{t}=-\infty$ and $\text{upper}_{t}=\mu_{t}+h_{t}$, with a time-varying mean $\mu_{t}$ and a bandwidth $h_{t}$. This band is calibrated on successful rollouts in $\mathcal{D}_{\text{eval-seen}}$. Under mild assumptions [59, 60], for any new successful rollout, $s_{t}<\mu_{t}+h_{t}$ holds for all $t=1,\dots,T$ with probability $1-\alpha$. Intuitively, this gives a guarantee that the false positive rate of the failure detector (a failure flag is raised at any time during a successful rollout) is at most $\alpha$. We use $\text{upper}_{t}$ as the failure flag threshold $\delta_{t}$, and more details about functional CP can be found in Appendix.
@@ 5.1 Evaluation Benchmarks
Real-robot Experiments: We deploy the $\pi_{0}$-FAST-DROID checkpoint [^1] https://github.com/Physical-Intelligence/openpi provided by [4, 5] on a Franka Emika Panda Robot. This checkpoint has been finetuned on the DROID dataset [33], and we do not further collect demonstrations or finetune the VLA model. We design 13 tasks and collect 30 successful and 30 failed rollouts for each task. The real-robot setup and example rollouts are visualized in Fig. 2. In experiments, 3 tasks out of 13 are randomly selected as unseen tasks.
Figure 3: Performance of functional CP by varying significant level $\alpha$ on $\mathcal{D}_{\text{eval-unseen}}$. The plots show the change of balanced accuracy (bal-acc) w.r.t. average detection time (T-det). Good failure detection methods should detect policy failures both accurately (high bal-acc) and proactively (lower T-det), and thus place curves towards the top left in each plot. We can see in all simulation experiments, the proposed SAFE-LSTM and SAFE-MLP perform better or on par with the best baselines. Note that baselines in gray require multiple action samples.
@@ 5.2 Uncertainty Quantification Baselines
Token uncertainty-based methods aggregate the predictive uncertainty from each generated token. These methods are efficient, as they only require a single forward inference. Given the generated tokens $\mathbf{W}_{t}=[\mathbf{w}_{t}^{1},\dots,\mathbf{w}_{t}^{m}]$, we denote the probability of sampling the token $\mathbf{w}_{t}^{i}$ as $p_{i}$ and the entropy over the distribution of the $i^{\text{th}}$ token as $H_{i}$. We adopt the token-based uncertainty estimation methods proposed in [23] as follows:
Token max prob.: $\max_{i}(-\log p_{i})$; Token avg prob.: $-\frac{1}{m}\sum_{i}\log p_{i}$;
@@
5.3 Failure Detection Baselines
Embedding Distance: We compare to baselines that directly use the distances in the feature space as failure scores. Specifically, instead of training a neural network, all VLA embeddings from $\mathcal{D}_{\text{train}}$ are stored in the two feature sets, $E_{\text{succ}}$ and $E_{\text{fail}}$, containing all VLA embeddings from successful and failed rollouts respectively. During evaluation on $\mathcal{D}_{\text{eval-seen}}$ and $\mathcal{D}_{\text{eval-unseen}}$, failure scores are computed as $s_{t}=d(\mathbf{e}_{t},E_{\text{succ}})-d(\mathbf{e}_{t},E_{\text{fail}})$, where $d(\cdot,\cdot)$ measures the distance between a single vector and a set of vectors. Intuitively, if $\mathbf{e}_{t}$ is far from $E_{\text{succ}}$ and close to $E_{\text{fail}}$, it’s more likely to fail. Following [17, 10], we ablate different types of distance, including Mahalanobis distance, and Euclidean and Cosine distance averaged over $k$-Nearest Neighbors of $\mathbf{e}_{t}$. We also compare to the PCA-KMeans distance measure from [9].
Learned OOD Detector: We adopt RND [39] and LogpZO [8], which are shown to be the best-performing OOD detection-based failure detectors by [8]. Both methods use a neural network $f^{\text{OOD}}(\cdot)$ to model the embedding distribution from successful rollouts and return an OOD score for a new embedding. We adapt them to learn from both successful and failed rollouts by training two models, $f^{\text{OOD}}_{\text{succ}}(\cdot)$ and $f^{\text{OOD}}_{\text{fail}}(\cdot)$, on $E_{\text{succ}}$ and $E_{\text{fail}}$ respectively. Similar to embedding distance baselines, the failure score is computed as $s_{t}=f^{\text{OOD}}_{\text{succ}}(\mathbf{e}_{t})-f^{\text{OOD}}_{\text{fail}}(\mathbf{e}_{t})$.
[…]
@@ 5.4 Evaluation Protocol
We consider two types of evaluation. The first type evaluates how well $s_{t}$ separates the successful and failed rollouts across all possible selections of $\delta_{t}$. Following the evaluation protocol widely adopted in the LLM UQ literature [49, 53, 23, 65], we use the area under ROC curve (ROC-AUC) metric. Furthermore, because a failure flag is raised whenever $s_{t}$ exceeds $\delta_{t}$, a successful rollout (ground truth negative) becomes a false positive whenever $s_{t}>\delta_{t}$, and remains a true negative only if $s_{t}\leq\delta_{t}$ for all time. Therefore, we consider the max-so-far score $\bar{s}_{t}=\max_{\tau=1,\dots,t}s_{\tau}$ and compute the ROC-AUC metric based on $\bar{s}_{T}$, the maximum failure score throughout the entire rollout.
[…]
@@ 6 Results
Figure 4: Failure scores predicted by SAFE-LSTM and the corresponding camera observations. The green areas show the functional CP band $C_{\alpha}$. Once failure scores exceed $C_{\alpha}$, a failure flag is raised. In (a), $\pi_{0}$-FAST policy misses the insertion, and its actions become unstable after that. In (b) and (c), OpenVLA and $\pi_{0}^{*}$ miss the grasp but still proceed to the placing action, causing a failure detection. Note that these tasks are not seen when training SAFE-LSTM.
@@ 6.1 How well do failure detectors distinguish failures from successes?
In Table 1 and Fig. 5 (a), we report the ROC-AUC metric based on $\bar{s}_{T}$, in simulation and real-world experiments, respectively. With a higher ROC-AUC metric, a failure detector achieves higher accuracy averaged over all possible thresholds. The tables show that Token Unc. methods have poor performance, which is aligned with findings in the LLM literature [24, 23]. On the other hand, the Sample Consist. and STAC [17] methods, which require multiple action samples, perform better and even achieve the best performance on unseen tasks in $\pi_{0}$-FAST LIBERO (STAC) and $\pi_{0}$ LIBERO (Gripper total var.). However, as these methods require multiple action samples, they cause significant overhead for VLA models and thus are not applicable to real robots. Embed. Distr. methods perform well, achieving the best performance in two simulation benchmarks ($\pi_{0}$ and $\pi_{0}$-FAST) and are the second best in the real world. This demonstrates that a VLA’s internal features are informative about task execution success/failure. The proposed SAFE methods perform better or on par with the best baselines, consistently in all settings. Averaged across simulation benchmarks, SAFE-MLP and SAFE-LSTM have similar performance, both outperforming the best baseline by 4-5% on unseen tasks, while still achieving the best performance on seen tasks. On the real robot, SAFE-MLP achieves the best performance and SAFE-LSTM performs closely with the best baseline (Euclid. $k$-NN). Comparing SAFE with Embed. Distr. methods, we attribute the success of SAFE to its stronger ability to extract high-level abstract information from raw feature vectors through learned neural networks.
Task Split Seen Unseen Max prob. 53.74 48.59 Avg prob. 51.60 47.30 Max entropy 59.23 53.50 Avg entropy 50.67 46.08 Mahala. dist. 75.54 53.93 Euclid. $k$-NN 80.35 60.27 Cosine. $k$-NN 80.23 59.51 PCA-KMeans 49.98 51.03 RND 62.00 45.83 LogpZO 64.43 52.24 STAC-Single 45.24 38.01 SAFE-LSTM 77.27 58.70 SAFE-MLP 86.76 64.16 (a) Failure Detection ROC-AUC
Figure 5: Quantitative results (a) and qualitative examples (b-e) from SAFE-MLP in the real world experiments. The ROC-AUC results are averaged over 5 random seeds with different task splits.
@@
6.2 How do detection accuracy and detection time trade off using functional CP?
@@
In Fig. 3, we use $\mathcal{D}_{\text{eval-seen}}$ to calibrate functional CP band $C_{\alpha}$ and evaluate on $\mathcal{D}_{\text{eval-unseen}}$. By varying the user-specified $\alpha$, we can adjust the conservativeness of the failure detectors and obtain a trade-off between accuracy (bal-acc) and detection time (T-det). To be a good failure detector, it should detect failures both accurately (higher bal-acc) and promptly (lower T-det), and thus have the curve rise toward the top-left corner in the plots of Fig. 3. As we can see from Fig. 3, the proposed SAFE-MLP and SAFE-LSTM perform the best on OpenVLA+LIBERO and $\pi_{0}$+SimplerEnv benchmarks, and on par with the best baseline on the other two benchmarks. We also manually annotate the ground truth (GT) failure timesteps (when a human thinks that failure happens or intervention is needed) for failed rollouts, and plot them as blue vertical lines in Fig. 3. Comparing SAFE’s performance with the GT fail time, we can see that SAFE can detect failures with high accuracy in the early stages of rollouts and potentially before the failure happens. This early detection allows early intervention for policy failures before they get stuck in execution or cause harm to the real-world environment.
@@
6.3 What are detected failure modes and are they aligned with human intuition?
@@
In Fig. 4 and Fig. 5(b-d), we visualize rollouts with the failure scores detected by SAFE. Fig. 4 demonstrates common failure modes in simulation, including imprecise insertion, unstable actions and missed grasps. Two successful rollouts on the real robot are shown in Fig. 5(b-c), where failure scores stop increasing after task completion. For the failed rollouts, the failure flag is raised after the policy is frozen (Fig. 5d) or the object slips out of the gripper (Fig. 5e). This aligns well with human intuition. Please refer to Appendix for video illustrations.

### v2

@@ 4.2 Failure Detection by Feature Probing
Table 1: Failure detection results on simulation benchmarks, measured by area under ROC (ROC-AUC). “-” indicates that the failure detection method does not apply. Entries with gray background indicate the failure detection methods that sample 10 actions per inference timestep, while others use only 1 action. The first and second best-performing methods are colored in red and orange, respectively. Results are averaged over 3 random seeds with different splits of seen and unseen tasks. SAFE achieves the highest averaged ROC-AUC over all simulation benchmarks.
[…]
@@ 4.3 Threshold Selection by Conformal Prediction
When the predicted failure score $s_{t}$ exceeds the time-varying threshold $\delta_{t}$, we raise a failure flag. To determine $\delta_{t}$ in a principled way, we adopt the functional conformal prediction (CP) framework [21]. Functional CP constructs a time-varying prediction band by leveraging the distribution of $s_{t}$ observed in successful rollouts within a calibration set. Under the exchangeability assumption [60] and given a user-specified significance level $\alpha$, the CP band guarantees that, for a new successful rollout, its $s_{t}$ will lie within this band at all times $t$ with probability $1-\alpha$. Conversely, if the score of a test rollout exceeds the band at time $t$, we can declare a failure with nominal confidence $1-\alpha$.
Formally, given a time series of any scalar score $s_{t}$ and a user-specified significance level $\alpha\in(0,1)$, functional CP gives a distribution-free prediction band $C_{\alpha}$. Following Xu et al. [8], we adopt the one-sided time-varying CP band formulation, where $C_{\alpha}$ is a set of intervals $\{[\text{lower}_{t},\text{upper}_{t}]:t=1,\dots,T\}$, where $\text{lower}_{t}=$-$\infty$ and $\text{upper}_{t}=\mu_{t}+h_{t}$, with a time-varying mean $\mu_{t}$ and a bandwidth $h_{t}$. This band is calibrated on successful rollouts in $\mathcal{D}_{\text{eval-seen}}$. Under mild assumptions [61, 62], for any new successful rollout, $s_{t}<\mu_{t}+h_{t}$ holds for all $t=1,\dots,T$ with probability $1-\alpha$. Intuitively, this gives a guarantee that the false positive rate of the failure detector (a failure flag is raised at any time during a successful rollout) is at most $\alpha$. We use $\text{upper}_{t}$ as the failure flag threshold $\delta_{t}$, and more details about functional CP can be found in Appendix.
@@ 5.1 Evaluation Benchmarks
Real-world Franka Experiments: We deploy the $\pi_{0}$-FAST-DROID checkpoint [4, 5] [^1] https://github.com/Physical-Intelligence/openpi on a Franka Emika Panda Robot. This checkpoint has been finetuned on the DROID dataset [32], and we do not further collect demonstrations or finetune the VLA model. We design 13 tasks and collect 30 successful and 30 failed rollouts for each task. The real-robot setup and example rollouts are visualized in Fig.˜3. In experiments, 3 tasks out of 13 are randomly selected as unseen tasks.
Real-world WidowX Experiments: We also deploy the OpenVLA model pretrained on the “Open-X Magic Soup++” dataset [2] on a WidowX robot manipulator in our lab. With this setup, we collected a total of 532 rollouts on the 8 lifting and pick-and-place tasks, including 244 successful and 288 failed rollouts. In this experiment, 2 tasks out of 8 are randomly selected as unseen tasks.
@@ 5.2 Uncertainty Quantification Baselines
Token uncertainty-based methods aggregate the predictive uncertainty from each generated token. These methods are efficient, as they only require a single forward inference. Given the generated tokens $\mathbf{W}_{t}=[\mathbf{w}_{t}^{1},\dots,\mathbf{w}_{t}^{m}]$, we denote the probability of sampling the token $\mathbf{w}_{t}^{i}$ as $p_{i}$ and the entropy over the distribution of the $i^{\text{th}}$ token as $H_{i}$. We adopt the token-based uncertainty estimation methods used by Huang et al. [22] as follows:
Token max prob.: $\max_{i}(-\log p_{i})$; Token avg prob.: $\frac{1}{m}\sum_{i}\log p_{i}$;
@@
Furthermore, inspired by semantic entropy [23], we define cluster entropy as entropy(cluster($\mathbf{A}_{t}^{k}\}_{k=1,\dots,K}$)), where cluster()($\cdot$) generates an integer set, containing $k$ cluster labels for the $k$ actions and entropy()($\cdot$) measures the entropy of the integer set. UQ methods based on sample consistency are shown to perform better for LLM [23, 22], but they necessitate multiple inferences, which may not be practical for large VLAs that control robots in real time.
Figure 4: In all simulation experiments, the proposed SAFE-LSTM and SAFE-MLP perform better than or on par with the best baselines. The plots show the variation of balanced accuracy (bal-acc) with respect to average detection time (T-det) on $\mathcal{D}_{\text{eval-unseen}}$, under different significance levels $\alpha$ used for functional CP. Good failure detection methods should detect policy failures both accurately (high bal-acc) and proactively (lower T-det), and thus place curves towards the top left in each plot. Note that baselines in gray require multiple action samples.
5.3 Failure Detection Baselines
Embedding Distance: We compare to baselines that directly use the distances in the feature space as failure scores. Specifically, instead of training a neural network, all VLA embeddings from $\mathcal{D}_{\text{train}}$ are stored in the two feature sets, $E_{\text{succ}}$ and $E_{\text{fail}}$, containing all VLA embeddings from successful and failed rollouts, respectively. During evaluation on $\mathcal{D}_{\text{eval-seen}}$ and $\mathcal{D}_{\text{eval-unseen}}$, failure scores are computed as $s_{t}=d(\mathbf{e}_{t},E_{\text{succ}})-d(\mathbf{e}_{t},E_{\text{fail}})$, where $d(\cdot,\cdot)$ measures the distance between a single vector and a set of vectors. Intuitively, if $\mathbf{e}_{t}$ is far from $E_{\text{succ}}$ and close to $E_{\text{fail}}$, it’s more likely to fail. Following recent works [18, 10, 66], we ablate different types of distance, including Mahalanobis distance, and Euclidean and Cosine distance averaged over $k$-Nearest Neighbors of $\mathbf{e}_{t}$. We also compare to the PCA-KMeans distance measure from Liu et al. [9].
Learned OOD Detector: Following Xu et al. [8], we adopt LogpZO, the best-performing method, and RND [39], a strong baseline, for OOD detection–based failure detectors. Both methods use a neural network $f^{\text{OOD}}(\cdot)$ to model the embedding distribution from successful rollouts and return an OOD score for a new embedding. We adapt them to learn from both successful and failed rollouts by training two models, $f^{\text{OOD}}_{\text{succ}}(\cdot)$ and $f^{\text{OOD}}_{\text{fail}}(\cdot)$, on $E_{\text{succ}}$ and $E_{\text{fail}}$ respectively. Similar to embedding distance baselines, the failure score is computed as $s_{t}=f^{\text{OOD}}_{\text{succ}}(\mathbf{e}_{t})-f^{\text{OOD}}_{\text{fail}}(\mathbf{e}_{t})$.
@@ 5.3 Failure Detection Baselines
Figure 5: Failures detected by SAFE-LSTM align well with the actual robot failures, as shown in the corresponding camera observations from simulation experiments. The blue-shaded areas show the functional CP band $C_{\alpha}$. Once failure scores exceed $C_{\alpha}$, a failure flag is raised. In (a), the $\pi_{0}$-FAST policy misses the insertion, and its actions become unstable after that. In (b) and (c), OpenVLA and $\pi_{0}^{*}$ miss the grasp but still proceed to the placing action, causing a failure detection. Note that these tasks are not seen when training SAFE-LSTM.
@@ 5.4 Evaluation Protocol
We consider two types of evaluation. The first type evaluates how well $s_{t}$ separates the successful and failed rollouts across all possible selections of $\delta_{t}$. Following the evaluation protocol widely adopted in the LLM UQ literature [49, 53, 22, 68], we use the area under the ROC curve (ROC-AUC) metric. Furthermore, because a failure flag is raised whenever $s_{t}$ exceeds $\delta_{t}$, a successful rollout (ground truth negative) becomes a false positive whenever $s_{t}>\delta_{t}$, and remains a true negative only if $s_{t}\leq\delta_{t}$ for all time. Therefore, we consider the max-so-far score $\bar{s}_{t}=\max_{\tau=1,\dots,t}s_{\tau}$ and compute the ROC-AUC metric based on $\bar{s}_{T}$, the maximum failure score throughout the entire rollout.
The second type of evaluation utilizes $\delta_{t}=\text{upper}_{t}$ calibrated by functional CP in Section˜4.3. By setting the significance level $\alpha$, we get a decisive positive/negative detection for each rollout. Following related works [18, 8], we consider the following metrics: true positive rate (TPR), false positive rate (FPR), balanced accuracy (bal-acc), and averaged detection time (T-det), where $\text{Bal-Acc}=\frac{\text{TPR+TNR}}{2}$. T-det is the relative timestep where $s_{t}>\delta_{t}$ for the first time (if $s_{t}$ never exceeds $\delta_{t}$, T-det becomes 1), averaged over all ground truth failed rollouts.
@@ 6.1 How well do failure detectors distinguish failures from successes?
In Table˜1 and Fig.˜6 (a), we report the ROC-AUC metric based on $\bar{s}_{T}$, in simulation and real-world experiments, respectively. With a higher ROC-AUC metric, a failure detector achieves higher accuracy averaged over all possible thresholds. The tables show that Token Unc. methods have poor performance, which is aligned with findings in the LLM literature [23, 22]. On the other hand, the Sample Consist. and STAC [18] methods, which require multiple action samples, perform better and even achieve the best performance on unseen tasks in $\pi_{0}$-FAST LIBERO (STAC) and $\pi_{0}$ LIBERO (Gripper total var.). However, as these methods require multiple action samples, they cause significant overhead for VLA models and thus are not currently practical for real robots. Embed. Distr. methods perform well, achieving the best performance in two simulation benchmarks ($\pi_{0}$ and $\pi_{0}$-FAST) and are the second best in the real world. This demonstrates that a VLA’s internal features are informative about task execution success/failure. The proposed SAFE methods perform better or on par with the best baselines, consistently in all settings. Averaged across simulation benchmarks, SAFE-MLP and SAFE-LSTM have similar performance, both outperforming the best baseline by 4-5% on unseen tasks, while still achieving the best performance on seen tasks. For the real-robot experiments, on both $\pi_{0}$-FAST+Franka and OpenVLA+WidowX rollouts, SAFE-MLP achieves the best performance and SAFE-LSTM performs closely with the best baseline (Mahala. dist. and Euclid. $k$-NN). Comparing SAFE with Embed. Distr. methods, we attribute the success of SAFE to its stronger ability to extract high-level abstract information from raw feature vectors through learned neural networks.
@@
6.2 What is the trade-off between detection accuracy and time using functional CP?
@@
In Fig.˜4, we use $\mathcal{D}_{\text{eval-seen}}$ to calibrate the functional CP band $C_{\alpha}$ and evaluate on $\mathcal{D}_{\text{eval-unseen}}$. By varying the user-specified $\alpha$, we can adjust the conservativeness of the failure detectors and obtain a trade-off between accuracy (bal-acc) and detection time (T-det). A good failure detector should detect failures both accurately (higher bal-acc) and promptly (lower T-det), and thus have the curve rise toward the top-left corner in the plots of Fig.˜4. As we can see from Fig.˜4, the proposed SAFE-MLP and SAFE-LSTM perform the best on OpenVLA+LIBERO and $\pi_{0}$+SimplerEnv benchmarks, and are on par with the best baseline on the other two benchmarks. We also manually annotate the ground truth (GT) failure timesteps (when a human thinks that failure happens or intervention is needed) for failed rollouts, and plot them as blue vertical lines in Fig.˜4. Comparing SAFE’s performance with the GT fail time, we can see that SAFE can detect failures with high accuracy in the early stages of rollouts and potentially before the failure happens. This early detection allows early intervention for policy failures before they get stuck in execution or cause harm to the real-world environment.
$\pi_{0}$-FAST Franka OpenVLA WidowX Method Seen Unseen Seen Unseen Max prob. 53.74 48.59 50.77 54.25 Avg prob. 51.60 47.30 48.94 44.36 Max entropy 59.23 53.50 51.88 49.19 Avg entropy 50.67 46.08 47.72 53.84 Mahala. dist. 75.54 53.93 82.37 70.00 Euclid. $k$-NN 80.35 60.27 72.01 53.64 Cosine $k$-NN 80.23 59.51 74.76 65.88 PCA-KMeans 49.98 51.03 75.62 47.22 RND 62.00 45.83 66.68 47.67 LogpZO 64.43 52.24 62.94 51.32 STAC-Single 45.24 38.01 SAFE-LSTM 77.27 58.70 84.29 71.80 SAFE-MLP 86.76 64.16 89.11 88.42 (a) Failure Detection ROC-AUC
Figure 6: SAFE-MLP achieves the best failure detection performance in real-world experiments with both $\pi_{0}$-FAST Franka and OpenVLA WidowX. Plot (a) presents quantitative results, while (b–e) show qualitative examples from SAFE-MLP on the real robot. ROC-AUC values are averaged over five random seeds with different task splits.
@@
6.3 What failure modes are detected, and do they align with human judgment?
@@
In Fig.˜5 and Fig.˜6(b-d), we visualize rollouts with the failure scores detected by SAFE. Fig.˜5 demonstrates common failure modes in simulation, including imprecise insertion, oscillatory motions, and missed grasps. Two successful rollouts on the real robot are shown in Fig.˜6(b-c), where failure scores stop increasing after task completion. For the failed rollouts, the failure flag is raised after the policy is frozen (Fig.˜6d) or the object slips out of the gripper (Fig.˜6e). This aligns well with human intuition. Please refer to Appendix for video illustrations.
6.4 How efficient and practical it is to deploy SAFE?
SAFE uses a 1-2 layer MLP or LSTM and poses a minimal (less than 1%) computational overhead at runtime. For example, SAFE-LSTM contains 2.3 million parameters and introduces an additional 0.73 ms of inference time. This is negligible compared to large VLA models. For instance, pi0 has 3.3 billion parameters and an inference time of 149 ms. SAFE only requires access to the latent features of VLA models and is compatible with any white-box robot policies based on neural networks. However, SAFE does require deploying the policy and collecting successful and failed rollouts to train the failure detector before it can detect failures.